OpenStudy (anonymous):

Find the pattern and the next number in the pattern. 1,8,27,64,125,.........

OpenStudy (agent47):

1^3 2^3 3^3 4^3 5^3 next = 6^3

OpenStudy (anonymous):

1^3 = 1 x 1 x 1 = 1; 2^3 = 2 x 2 x 2 = 8; 3^3 = 3 x 3 x 3 = 27; 4^3 = 4 x 4 x 4 = 64; 5^3 = 5 x 5 x 5 = 125